OpenStudy (anonymous):

how do you solve this?? The sum of three consecutive numbers is 132

OpenStudy (anonymous):

here is how you solve this, you get the number and devide it by the amount of consecutive numbers. then you use the number that it came out to as the median and count behind and in front of that number to get the numbers your looking for example: 132/3=44 43+44+45=132 hope this helps

OpenStudy (solomonzelman):

3 consecutive numbers can be written as x , (x+1) , (x+2) their sum would therefore be x+(x+1)+(x+2) we know that their sum is equal to 132, so x+(x+1)+(x+2) =132 WAY 1: x+(x+1)+(x+2) =132 x+x+1+x+2 =132 3 x+3 =132 3x=129 x=43 So x (x+1) (x+2) would be 43, 44, 45 WAY 2: 3 consecutive number add up to 132, so their average (or the middle number of the three consecutive numbers) (and out of 3 terms) is equal to 132. a= 132/3 a=44 knowing that the middle number is 44, the numbers are 43, 44, 45

OpenStudy (tkhunny):

Don't get caught in one way of thinking. Rather than define the least of the three as x, try defining the middle value as x (x-1) + x + (x+1) = 3x = 132 x = 44 and we have 44 - 1 = 43, 44, 44 + 1 = 45 Unique answers don't care how you find them. Try to find ways that make your life easier.